# The development of a high granular crystal calorimeter prototype of VLAST

### v1

@@ III Construction of the HEIC-Cube prototype
Figure 7: The schematic plot of the high granular cubic crystal calorimeter prototype.
The mechanical design of the prototype is carefully considered to ensure the precise positioning and stability of the BGO crystals and readout electronics. Fig. 8 shows the box of the sensitive layer made of carbon fiber, a lightweight and strong material, which consists of a 5 × 5 array of arranged cells. Each cell measures 31.6 mm × 31.6 mm to accommodate a BGO crystal placement with a small tolerance for positioning and thermal expansion. The intercellular partitions have a uniform thickness of 1 mm, while the outermost frames exhibit a 2 mm thickness for additional structural reinforcement. Fig. 8 demonstrates the assembled configuration, where BGO crystals are precisely positioned within their respective cell and covered with attenuation filters above their optical windows. The filters are crucial for the dynamic range extension of the readout system, as they attenuate the light reaching one of the two APDs coupled to each crystal, which allows the system to measure a wider range of energies without saturation. The crystals are securely fixed in place using a black shock-absorbing adhesive (DOWSIL SE 9186 L Black Sealant). This adhesive provides both mechanical stability and optical isolation. Additionally, a carbon fiber honeycomb structure is placed above the filters to maintain the required air gap.
The prototype readout electronic board is shown in Fig. 8, housing the readout for 25 crystals arranged in a horizontal layer. APDs have been soldered onto the PCB for efficient signal transfer and mechanical stability, and each crystal unit interfaces with 2 APDs. They share the same voltage supplied by a Nuclear Instrumentation Module (NIM) power source, and a compact high voltage low dropout regulator (LDO) will be installed at the top right-hand corner of PAM during subsequent assembly. The PAM board is securely buckled onto the carbon fiber box according to locating bolts, while the corresponding ADM modules are connected and firmly fixed using screws, thereby constituting a complete sensitive layer as depicted in Fig. 8. The complete prototype configuration integrates 10 identical layers in vertical stacking and a corresponding back-end DCM.
@@
In summary, the cubic crystal calorimeter prototype comprises 10 detection layers, with each layer consisting of 25 BGO crystals with a side length of 30 mm. Each crystal is coupled with two APDs, generating 4 distinct amplitude output signals. This design results in a system-wide configuration of 250 crystals and 960 active readout channels. Accordingly, the detector’s electromagnetic shower containment capability is characterized by 3.3 Molière radii in the transverse dimension and 26.8 radiation lengths in the longitudinal dimension when particles hit the calorimeter along the central axis. These parameters ensure that the prototype is well-suited for precise electron energy measurement in the GeV range. The prototype provides valuable experimental data for validating the performance of the dual-APD, dual-gain readout scheme and for characterizing the calorimeter’s response to electromagnetic showers, paving the way for the development of the full-scale VLAST calorimeter.
@@ IV Performance of the prototype system
Figure 9: Overview photograph of the cubic calorimeter prototype when conducting cosmic ray test.
To further validate the calorimeter prototype and assess its performance under real-world conditions, an experimental setup was established for ground-based cosmic ray measurements. This testing involves exposing the prototype to the natural flux of cosmic rays, which serve as a readily available source of high-energy particles, providing an opportunity to evaluate its response. Fig. 9 depicts the experimental setup for this cosmic ray test.
@@
The data acquisition system employed in this experiment utilizes an over-threshold triggering mode, which means that data is recorded only when the signal in a readout channel exceeds a predefined threshold. This threshold is set to 20 times the sigma value relative to its respective pedestal average for each channel. The pedestal represents the baseline signal level of each channel when no particle is interacting with the detector, while its width reflects the level of noise. Setting the threshold at 20 times the sigma value helps to discriminate real signals from electronic noise fluctuations.
@@
Some preliminary analysis results are shown here from the routine cosmic ray test. The offline data processing involves several steps to extract meaningful information from the raw data. First, the binary files are unpacked to retrieve the 512 sampling points datasets for each channel. These sampling points represent the digitized signal waveform over time. The average value of the first 128 sampling points is calculated and used as the pedestal value for that channel. The maximum code value among all 512 points is identified as the channel’s peak value, representing the maximum amplitude of the signal. The effective amplitude of each channel is then determined by subtracting the pedestal value from the peak value. This procedure is applied consistently across all daily data packets to ensure uniformity in the analysis.
@@
Further analysis reveals significant variations in the mean pedestal values between different channels, as illustrated in subsequent figures ( 10 and 10) showing the mean values derived from Gaussian fitting of high-gain and low-gain channels during a one-day test. These variations can be attributed to several factors, including non-uniformity of temperature across the BGO crystals and APDs at different spatial locations, disparities in gain coefficients among APDs despite the common bias high voltage, and potential slight light leakage effects. Temperature variations can affect the performance of both the BGO crystals and the APDs, leading to changes in their response. Variations in the gain coefficients of the APDs can arise from manufacturing tolerances or differences in their operating conditions and recommended setup. Light leakage can introduce unwanted background signals, affecting the pedestal values. These factors exert various degrees of influence on the signal generation and amplification processes, thereby endowing each channel’s pedestal with distinct characteristics.
Figure 10: (a) An example of pedestal distribution in HH12 channel Layer 7. (b) An example of pedestal distribution in HL12 channel Layer 7. (c) and (d) 2 show statistical results of all the mean value of pedestals in high-gain channels and low-gain channels from 1 day’s test.
To determine the gain coefficients between the high-gain and low-gain channels, a linear fitting analysis is performed. The amplitudes relationship of the two output signals from the same APD without attenuation filter coverage, are analyzed before and after amplification by the front-end electronics, as presented in Fig. 11. The linear region of the relationship between the two signals is fitted using a first-order polynomial function. The average slope value obtained from this fit is approximately 37.5, indicating that these gain coefficients are similar and maintain reasonable consistency with the original design specifications. Given that the energy deposition of muons in cosmic rays is generally relatively small, only the uncovered APDs with high-gain can produce detectable MIP signals. Therefore, it is challenging to calibrate the remaining two coefficients (HL-LH ratio and LH-LL ratio) using a single data file.
@@
A group of specific selection criteria are defined to identify MIP events within the dataset. A MIP event is defined as an event in which at least 7 out of the 10 detection layers output effective signals, with no more than 3 over-threshold channels per layer, and there should be at least 2 of the first 3 layers and 2 of the last 3 layers meet these criteria. This selection ensures that the identified events correspond to muons traversing the entire calorimeter. Fig. 12 depicts a representative event that satisfies these requirements. The selected MIP events are then analyzed channel by channel. The signal distribution for each channel is fitted using a Landau convoluted Gaussian function, which is commonly used to model the energy loss distribution of charged particles traversing a material. The Most Probable Value (MPV) obtained from the Landau distribution component represents its response to MIP in a given crystal. For example, the HH12L7 channel exhibits an MPV of approximately 370 ADC counts after pedestal subtraction, corresponding to an estimated input charge of about 18 fC for MIP signals.
@@
Figure 12: A typical event which are selected as MIPs.
Figure 13: (a) A typical MIPs distribution after subtracting its own pedestal in HH12 channel Layer 7. (b) A statistical result of all the most probable value of MIPs in high-gain channels from 1 day’s test.
The long-term stability of the calorimeter’s response was also investigated. The cosmic ray experiment has been running for two months, allowing for the observation of long-term trends. Some parameters, such as the pedestal and MPV of Landau component, showed gradually emerging regular patterns over time. Fig. 14 demonstrates that the MPV of individual channels varied within a relatively small range, while the pedestal values remained remarkably stable throughout the testing period. The temperature of the channels, shown as dotted lines in Fig. 14 with reference to the right axis, exhibits a significant correlation with the observed signal variations, further supporting the hypothesis that temperature fluctuations play a significant role in the observed channel-to-channel variations. This long-term stability is crucial for ensuring the reliable performance of the calorimeter over extended periods of operation. Consequently, more stringent temperature control requirements have been put forward for the satellite’s payload platform.

### v2

@@ III Construction of the HEIC-Cube prototype
Figure 7: The schematic plot of the high granular crystal calorimeter prototype.
Prior to the construction of the calorimeter, a pairing process is carried out based on the fluorescence output of each BGO crystal and the gains of its corresponding APDs. This ensures a better consistency in the response of each detection unit to MIPs. These matched crystal-APDs pairs are then installed in corresponding positions within the same layer to maintain similar performance characteristics. The mechanical design of the prototype is carefully considered to ensure the precise positioning and stability of the BGO crystals and readout electronics. Fig. 8 shows the box of the sensitive layer made of carbon fiber, a lightweight and strong material, which consists of a 5 × 5 array of arranged cells. Each cell measures 31.6 mm × 31.6 mm to accommodate a BGO crystal placement with a small tolerance for positioning and thermal expansion. The intercellular partitions have a uniform thickness of 1 mm, while the outermost frames exhibit a 2 mm thickness for additional structural reinforcement. Fig. 8 demonstrates the assembled configuration, where BGO crystals are precisely positioned within their respective cell and covered with attenuation filters above their optical windows. The filters are crucial for the dynamic range extension of the readout system, as they attenuate the light reaching one of the two APDs coupled to each crystal, which allows the system to measure a wider range of energies without saturation. The crystals are securely fixed in place using a black shock-absorbing adhesive (DOWSIL SE 9186 L Black Sealant). This adhesive provides both mechanical stability and optical isolation. Additionally, a carbon fiber honeycomb structure is placed above the filters to maintain the required air gap.
The prototype readout electronic board is presented in Fig. 8, housing the readout for 25 crystals arranged in a horizontal layer. APDs have been soldered onto the PCB for efficient signal transfer and mechanical stability, and each crystal unit interfaces with 2 APDs. They share the same voltage supplied by a Nuclear Instrumentation Module (NIM) power source, and a compact high voltage low dropout regulator (LDO) will be installed at the top right-hand corner of PAM during subsequent assembly. The consistency of the high-gain and low-gain coefficients in the readout electronics is relatively straightforward to control. The PAM board is securely buckled onto the carbon fiber box according to locating bolts, while the corresponding ADM modules are connected and firmly fixed using screws, thereby constituting a complete sensitive layer as depicted in Fig. 8. The complete prototype configuration integrates 10 identical layers in vertical stacking and a corresponding back-end DCM (Fig. 8). Following assembly, each detection unit must still undergo calibration using MIPs. Their responses are subsequently normalized based on the respective MIP signal amplitude, thereby harmonizing the output across all units and enhancing the overall uniformity and measurement reliability of the detector system.
@@
In summary, the high granular crystal calorimeter prototype comprises 10 detection layers, with each layer consisting of 25 BGO crystals with a side length of 30 mm. Each crystal is coupled with two APDs, generating 4 distinct amplitude output signals. This design results in a system-wide configuration of 250 crystals and 960 active readout channels. Accordingly, the detector’s electromagnetic shower containment capability is characterized by 3.3 Molière radii in the transverse dimension and 26.8 radiation lengths in the longitudinal dimension when particles hit the calorimeter along the central axis. These parameters ensure that the prototype is well-suited for precise electron energy measurement in the GeV range. The prototype provides valuable experimental data for validating the performance of the dual-APD, dual-gain readout scheme and for characterizing the calorimeter’s response to electromagnetic showers, paving the way for the development of the full-scale VLAST calorimeter.
@@ IV Performance of the prototype system
Figure 9: Overview photograph of the calorimeter prototype when conducting cosmic ray test.
To further validate the calorimeter prototype and assess its performance under real-world conditions, an experimental setup was established for ground-based cosmic ray measurements. This testing involves exposing the prototype to the natural flux of cosmic rays, which serve as a readily available source of high-energy particles, providing an opportunity to evaluate its response. Fig. 9 demonstrates the experimental setup for this cosmic ray test.
@@
The data acquisition system employed in this experiment utilizes an over-threshold triggering mode, which means that data is recorded only when the signal in a readout channel exceeds a predefined threshold. This threshold is set to 20 times the sigma value relative to its respective pedestal average for each channel. The pedestal represents the baseline signal level of each channel when no particle is interacting with the detector, while its width reflects the level of noise. Setting the threshold at 20 times the sigma value (approximate 130 ADC values, equivalent to 0.3 MIP MPV) helps to discriminate real signals from electronic noise fluctuations.
@@
Some preliminary analysis results are presented here from the routine cosmic ray test. The offline data processing involves several steps to extract meaningful information from the raw data. First, the binary files are unpacked to retrieve the 512 sampling points datasets for each channel. These sampling points represent the digitized signal waveform over time. The average value of the first 128 sampling points is calculated and used as the pedestal value for that channel. The maximum code value among all 512 points is identified as the channel’s peak value, representing the maximum amplitude of the signal. The effective amplitude of each channel is then determined by subtracting the pedestal value from the peak value. This procedure is applied consistently across all daily data packets to ensure uniformity in the analysis.
@@
Further analysis reveals significant variations in the mean pedestal values between different channels, as depicted in subsequent figures ( 10 and 10) showing the mean values derived from Gaussian fitting of high-gain and low-gain channels during a one-day test. These variations can be attributed to several factors, including non-uniformity of temperature across the BGO crystals and APDs at different spatial locations, disparities in gain coefficients among APDs despite the common bias high voltage, and potential slight light leakage effects. Temperature variations can affect the performance of both the BGO crystals and the APDs, leading to changes in their response. Variations in the gain coefficients of the APDs can arise from manufacturing tolerances or differences in their operating conditions and recommended setup. Light leakage can introduce unwanted background signals, affecting the pedestal values. These factors exert various degrees of influence on the signal generation and amplification processes, thereby endowing each channel’s pedestal with distinct characteristics.
Figure 10: (a) An example of pedestal distribution in HH12 channel Layer 7. (b) An example of pedestal distribution in HL12 channel Layer 7. (c) and (d) demonstrate statistical results of all the mean value of pedestals in high-gain channels and low-gain channels from 1 day’s test. The conversion factor in the high gain channel is 20.6 ADC/fC, while the factor in the low gain channel is 0.53 ADC/fC, according to bib:58 .
To determine the gain coefficients between the high-gain and low-gain channels, a linear fitting analysis is performed. The amplitudes relationship of the two output signals from the same APD without attenuation filter coverage is presented in Fig. 11. The linear region of the relationship between the two signals is fitted using a first-order polynomial function. Fig. 11 collects and demonstrates the slope value after linear fitting, and their average is approximately 37.5, indicating that these gain coefficients are similar and maintain reasonable consistency with the original design specifications. Given that the energy deposition of muons in cosmic rays is generally relatively small, only the uncovered APDs with high-gain can produce detectable MIP signals. Therefore, it is challenging to calibrate the remaining two coefficients (HL-LH ratio and LH-LL ratio) using a single data file.
@@
A group of specific selection criteria are defined to identify MIP events within the dataset. A MIP event is defined as an event in which at least 7 out of the 10 detection layers output effective signals, with no more than 3 over-threshold channels per layer, and there should be at least 2 of the first 3 layers and 2 of the last 3 layers meet these criteria. This selection ensures that the identified events correspond to muons traversing the entire calorimeter. Fig. 12 depicts a representative event that satisfies these requirements. The selected MIP events are then analyzed channel by channel. The signal distribution for each channel is fitted using a Landau convoluted Gaussian function, which is commonly used to model the energy loss distribution of charged particles traversing a material. The Most Probable Value (MPV) obtained from the Landau distribution component represents its response to MIP in a given crystal. For example, the HH12L7 channel illustrated in Fig. 13 exhibits an MPV of approximately 370 ADC counts after pedestal subtraction, corresponding to an estimated input charge of about 18 fC for MIP signals.
@@
Figure 12: A typical event which is selected as a MIP incident event.
Figure 13: (a) A typical MIPs distribution after subtracting its own pedestal in HH12 channel Layer 7. (b) A statistical result of all the most probable value of MIPs in high-gain channels from 1 day’s test. The conversion factor in the high gain channel is 20.6 ADC/fC, according to bib:58 .
The long-term stability of the calorimeter’s response was also investigated. The cosmic ray experiment has been running for two months, allowing for the observation of long-term trends. Some parameters, such as the pedestal and MPV of Landau component, showed gradually emerging regular patterns over time. Fig. 14 demonstrates that the MPV of individual channels varied within a relatively small range, while the pedestal values remained remarkably stable throughout the testing period. The temperature of the channels, presented as dotted lines in Fig. 14 with reference to the right axis, exhibits a significant correlation with the observed signal variations, further supporting the hypothesis that temperature fluctuations play a significant role in the observed channel-to-channel variations. This long-term stability is crucial for ensuring the reliable performance of the calorimeter over extended periods of operation. Consequently, more stringent temperature control requirements have been put forward for the satellite’s payload platform.